during a 10 minute period, a weigth lifter lifts a 30 kg dumbbell 50 times, from the floor level to a height of 1.75 meters. what is the average power ( in horse power) done by the weight lifter


1
Expert's answer
2022-02-24T09:05:30-0500

Average power done by the weight lifter is:


"P=\frac{Fs}{t}=\frac{mghn}{t}" ,


where F - the lifting force,

s - covered distance,

h - height,

n - number of liftings.


Assuming that 1 horse power is 745.7 W


"P=\frac{mghn}{t}=\frac{30\cdot 9.8\cdot 1.75 \cdot 50}{600}=42.9\space W\approx 0.06 \space horse\space power" .